x+y-2z=-1
2x-y+z=-6
3x+2y-3z=-3

Answer by Fermat(136)

x+y-2z=-1 ----------(1)
2x-y+z=-6 ----------(2)
3x+2y-3z=-3 --------(3)
from (1), x = -y+2z-1
substitute for x = -y+2z-1 into (2) and (3),
2(-y+2z-1) - y + z = -6
3(-y+2z-1) + 2y - 3z = -3
multiplying out,
-2y + 4z - 2 - y + z = -6
-3y + 6z - 3 + 2y - 3z = -3
rearranging,
-3y + 5z = -4 -------(4)
-y + 3z = 0 -------(5)
from (5), y = 3z
substitute for y = 3z into (4),
-3(3z) + 5z = -4
-9z + 5z = -4
-4z = -4
z = 1
=====
y = 3z
y = 3
=====
x = -y+2z-1
x = -3 + 2 - 1
x = -2
======
Ans: x=-2, y=3, z=1
===================